An engineer weighs a sample of mercury (ρ =
13.6 × 103 kg/m3) and finds that the weight
of the sample is 5.6 N.
What is the sample’s volume?
1
Expert's answer
2011-02-14T05:17:23-0500
The mass of the sample is m = W/g = 5.6 [N]/9.8[N/kg] = 0.57 kg.
The volume is V = m/ρ = 0.57[kg]/13.6x103[kg/m3] = 4.2x10-5 m3.